Presentation on theme: "KICKOFF AS PRESENTED 6/17/11, SEE THIS WEEKS SCHEDULE FOP THE LATEST UPDATE Shutdown Shutdown ‘11."— Presentation transcript:
KICKOFF AS PRESENTED 6/17/11, SEE THIS WEEKS SCHEDULE FOP THE LATEST UPDATE Shutdown Shutdown ‘11
Why this meeting? Discuss any last minute issues that may impact the shutdown schedule. Show schedule for the beginning of the shutdown. Ensure critical access are made on time. Ensure areas opened in accordance with the work scheduled. Begin Critical and very long jobs ASAP.
General Schedule The 27 GeV/n CM run will end 0600hrs June 29 th. Power supply work June 29 th 0600-1000hrs. The RHIC and AGS cold snake warm up will begin immediately after PS work. AGS will run proton studies until June 30 th. NSRL will un until the end June 30 th. ODH 1 postings remain in place until July 6 th. Obey all postings, they will be modified ASAP
General schedule continued Booster, AGS and RHIC Berms and MDs will be opened sequentially unless given higher priority… Week of July 6 th has rigging scheduled to remove blocks from access gates (MD). They will be left opened as blocks are removed. Radiation area postings at RHIC will be removed on July 1 st. AGS will remain locked until July 7 th. Friday, July 1 st : AMMPS personnel to connect Westinghouse to AGS ring. Wednesday, July 6 th : Pulse tests with Westinghouse into AGS. Booster will be opened following LOTO and RA surveys. Other locked areas (fan houses etc.) will be opened starting July 1 st.
Major RHIC systems shutdown: June 29 th. July 1 st : STAR and PHENIX systems to idle, magnets off. CAS to rack out supplies as necessary. RMMPS LOTO for the shutdown. IR and corrector supplies secured. Begin service building summer mode settings. Temperature set-points will be somewhat higher, Begin power saving measures for all RHIC systems Secure/LOTO AtR supplies
June 29 th RHIC: Address any last minute cold issues (10am June 29 at the latest). RHIC warm up begins June 29 th 10am. HP surveys in dump areas. RHIC RA access begins. Secure/LOTO RHIC equipment. Perform HP surveys, open AtR. MD and escape tunnel accesses will be opened during the week of July 6 th. Riggers will remove PHENIX shielding first, then blocks around the ring.
Pre Injectors End game: Booster run for NSRL ends June 30 th. Booster Vacuum will be isolated from EBIS to allow for continued commissioning. ETB “Big Red Bend” will remain energized with postings. EBIS commissioning schedule still developing. Will continue for some time TBD. LINAC run protons for BLIP and AGS studies until June 30 th. BLIP and NSRL off in afternoon or early evening. AGS studies most likely run until midnight.
Tunnel Access and Timing On June 29 th access will similar to a maintenance day: Following Power supply work, Areas except for the dumps will be opened to RA. After HP surveys, all dump areas will be opened For all access, assume running ODH levels until otherwise posted Normally locked areas should be opened by the end of the week For access status consult the web or TV displays
LINAC and Booster Running until June 30 th. EBIS work will continue (beam to but not injected in the Booster). July 1 st, Booster berms will be opened after de- posting. NSRL tunnel areas will be opened in the evening of June 30 th after the last user and final work and documentation is completed. LINAC tunnel will be opened after a cool down period TBD based on HP surveys.
Power outages: Electrical outages for maintenance or other reasons will be scheduled around critical path work. Weekly outage lists will be displayed on CCTV and on the web. Full outage schedules will be available on the web.
Full schedules and timelines: Will be available on the web “This Week” Discussed at meetings Supervisors meeting (weekly) Shutdown Meetings Bi-weekly, in format similar to time meeting. Start Up meetings (daily) Will begin as the run approaches Defines access and activity for the day
Shutdown work Presently Reviewing Submitted Jobs. Major project work schedules are being determined Some work and priorities and schedules will be revisited at the RHIC retreat July 20-22. The following slides are an outline of some of the major efforts planned. A complete list will be generated and posted on line. LIST OF JOBS
Major work at a glance: Injectors Tandem will conduct a general maintenance period and run for outside users as necessary. LINAC Maintenance to LINAC tanks and systems. New beryllium windows for BLIP. Booster Foil replacement at C3 and in BtA with bake-outs EBIS work continues… AGS: Sextupole refurbish continues. G17 gamma quad replace (roof off of AGS?). I5 quad vacuum repair and reinstall. Main Magnet coil replacement? Access controls…
RHIC e-lens Installation of infrastructure: Tunnel penetrations. Cable installation. Buss installation. Initial stand installation. Modifications to 1010A and 1010 trailer. Cooling system install in 1010 buildings. Install cryogenic tap.
RHIC Projects: Major re-work in RF area: Prep for 56MHz system: Cryogenic tap, compressor, trays, power supply racks Redesign of 9 MHz, Bouncer modification Install new landau cavities, modify systems and commission. 28 MHz system and cavity modifications. 197 MHz system and cavity modifications.
Stochastic cooling Fabrication and rebuilds: Blue and Yellow horizontal kickers, Blue longitudinal kicker. Repairs: yellow longitudinal. Installations: Blue & Yellow pickups and H kickers, Blue Longitudinal kicker, water systems for all tanks. Upgrades: Drives, low level, hardware and cabling. Other work…
2 o’clock AnDY work: Install PHOBOS magnet in IR. Modify HCAL detectors. Install ECAL detectors. Move existing e-cal detectors. Beryllium pipe remove and re-install. Cryogenic installation tap for CeC. Bake outs…
STAR Shutdown dependant on GEM tracker installation. In that case STAR would be rolled out for the shutdown. Possible impact on RHIC schedule. A new or modified vacuum pipe for the IR would be needed. Otherwise, STAR work would be minimal.
PHENIX Installation of the FVTX Re-design and installation of the shielding for the PRC north and South tunnels. Other work….